Ok dmodule (a library for D-modules by Okutani) Manual

Edition : auto generated by oxgentexi on 30 January 2008

OpenXM.org

We need to load the files ‘gr’, ‘xm’, ‘ok_matrix.rr’, ‘ok_diff.rr’, ‘ok_diff.rr’ to call functions in this manual.

The document has not yet been written. Here are examples.

[344] odmodule_ch_ideal([x*dx+y*dy+a,dx^2+dy^2],[x,y]);
[x*dx+y*dy,dx^2+dy^2,y*dy*dx-x*dy^2,(x^2+y^2)*dy^2]
[348] odmodule_ch_ideal(odiff_op_appell4(a,b,c1,c2,[x,y]),[x,y]);
[-x*dx^2+y*dy^2,2*y*x*dy*dx+(y*x+y^2-y)*dy^2,
 (2*y^2-2*y)*dy^2*dx+(-y*x+3*y^2+y)*dy^3,
 2*y*x*dy^2*dx+(y*x^2+(-2*y^2-y)*x+y^3-y^2)*dy^3]

Index 2

Index

(Index is nonexistent) (Index is nonexistent)

Short Contents

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

Table of Contents

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2